# Computation of Robust Option Prices via Structured Multi-Marginal Martingale Optimal Transport

### v1

@@ 4.3 Summary of the full algorithm
We have now arrived at a form of the optimality conditions that allows for efficient computation of the optimal dual variables. We can therefore construct a coordinate dual ascent method by cyclically fixing all but one variable and optimise over the remaining variable by selecting it so that it satisfies the corresponding optimality condition. We then move on to the next variable until all variables have been optimised over; one such cycle defines one iteration in the method.
Each iteration starts with the update of the variables $u$ corresponding to the marginal constraints (22c). The closed-form formula for doing so is given by Theorem 4.1 as
$u_{t}\leftarrow\upmu_{t}\oslash P^{S}(\hat{\psi}_{t}\odot\psi_{t}),\quad t\in\mathcal{T}.$
From the above we note that is suffices to manipulate vectors in order to update $u$, something that enables fast computation.
We then proceed to updating the dual variables $\gamma$ representing the martingale constraints (22b). This time, no closed-form formula exists for $\gamma$ — the optimality condition thus has to be solved numerically. Theorem 4.1 provides a form of the equations that allows for efficiently doing so by application of Newton’s method. In order to do so, define a family of functions $\alpha_{t}:\mathbb{R}^{n_{S}n_{X}}\rightarrow\mathbb{R}^{n_{S}n_{X}}$ for $t\in[T]\backslash 0$ via
$\alpha_{t}(z):=\begin{cases}(K_{t}\odot\exp(\operatorname{diag}(z)\Delta/\varepsilon)\odot\Delta)(\psi_{t}\odot(\mathbf{1}_{n_{X}}\otimes u_{t})),&t\in\mathcal{T}\\ (K_{t}\odot\exp(\operatorname{diag}(z)\Delta/\varepsilon)\odot\Delta)\psi_{t},&t\notin\mathcal{T}.\end{cases}$
Then equation 31 can be written
$\begin{cases}\hat{\psi}_{t}\odot(\mathbf{1}_{n_{X}}\otimes u_{t})\odot\alpha_{t+1}(\gamma_{t})=\mathbf{0}_{n_{S}n_{X}},&t\in[T-1]\cap\mathcal{T}\\ \hat{\psi}_{t}\odot\alpha_{t+1}(\gamma_{t})=\mathbf{0}_{n_{S}n_{X}},&t\in[T-1]\backslash\mathcal{T}.\end{cases}$
@@
$\gamma_{t}^{(k+1)}\leftarrow\gamma_{t}^{(k)}-\theta^{(k)}\odot J_{\alpha_{t+1}}\big{(}\gamma_{t}^{(k)}\big{)}^{-1}\alpha_{t+1}\big{(}\gamma_{t}^{(k)}\big{)},\quad k=1,2,\dots,$ (37)
where $\theta^{(k)}\in\mathbb{R}^{n_{S}n_{X}}$ is a vector containing the respective step lengths to be used for each variable, such that $\theta^{(k)}(i)\in[0,1]$ for $i\in\{1,\dots,n_{S}n_{X}\}$. It is determined by vectorised line search. In the above, $J_{\alpha_{t+1}}$ denotes the Jacobian matrix associated with $\alpha_{t+1}$; note that $\alpha_{t+1}$ is a vector-valued function whose $i^{\text{th}}$ component is independent of any other components of $\gamma_{t}$ than $\gamma_{t}(i)$, $i\in\{1,\dots,n_{S}\}\times\{1,\dots,n_{X}\}$. There are no cross dependencies. Therefore, the Jacobian is a diagonal matrix and given by $J_{\alpha_{t}}(\gamma_{t-1})=\operatorname{diag}(\beta_{t}(\gamma_{t-1}))$, where $\beta_{t}:\mathbb{R}^{n_{S}n_{X}}\rightarrow\mathbb{R}^{n_{S}n_{X}}$ for $t\in[T]\backslash 0$ is defined as
$\beta_{t}(z):=\begin{cases}\frac{1}{\varepsilon}(K_{t}\odot\exp(\operatorname{diag}(z)\odot\Delta/\varepsilon)\odot\Delta\odot\Delta)(\psi_{t}\odot(\mathbf{1}_{n_{X}}\otimes u_{t})),&t\in\mathcal{T}\\ \frac{1}{\varepsilon}(K_{t}\odot\exp(\operatorname{diag}(z)\odot\Delta/\varepsilon)\odot\Delta\odot\Delta)\psi_{t},&t\notin\mathcal{T}.\end{cases}$
@@
$\gamma_{t}^{(k+1)}\leftarrow\gamma_{t}^{(k)}-\theta^{(k)}\odot\alpha_{t+1}\big{(}\gamma_{t}^{(k)}\big{)}\oslash\beta_{t+1}\big{(}\gamma_{t}^{(k)}\big{)},\quad k=1,2,\dots$
The inversion of the Jacobian and the matrix-vector multiplication used in (37) are thus avoided and replaced by the elementwise division of one vector with another.
Putting things together, we have completed the derivation of our algorithm — see Algorithm 2 for the full method. Convergence of the algorihtm can be shown given that there exists a feasible solution that is positive for all variables that are not directly constrained to be zero by the constraints (cf. [32, Theorem 4.1 and Proposition 1]). In this case, the variables constrained to be zeros can be removed (following [32]); convergence of the coordinate dual ascent then follows from [45, Theorem 2 and Section 5].
Given: matrices $\{K_{t}\}_{t=1}^{T}$ and $\Delta$, index set $\mathcal{T}$, vectors $\{\upmu_{t}\}_{t\in\mathcal{T}}$, scalar $\varepsilon>0$
Initialise: $u_{t}\leftarrow\mathbf{1}_{n_{S}}$ for $t\in[T]$
Initialise: $\gamma_{t}\leftarrow\mathbf{1}_{n_{S}n_{X}}$ for $t\in[T-1]$
Initialise: $G_{t}\leftarrow\exp(\operatorname{diag}(\gamma_{t-1})\Delta/\varepsilon)$ for $t\in[T]\backslash 0$
Initialise: $\hat{\psi}_{1},\psi_{T}\leftarrow\mathbf{1}_{n_{S}n_{X}}$
Initialise: $\psi_{t}\leftarrow(K_{t+1}\odot G_{t+1})(\psi_{t+1}\odot(\mathbf{1}_{n_{X}}\otimes u_{t+1}))$ for $t\in[T-1]$
@@
         $u_{0}\leftarrow\upmu_{0}\oslash P^{S}(\psi_{0})$
@@
         $\hat{\psi}_{t}\leftarrow(K_{t}\odot G_{t})^{\top}(\hat{\psi}_{t-1}\odot(\mathbf{1}_{n_{X}}\otimes u_{t-1}))$
@@
             $u_{t}\leftarrow\upmu_{t}\oslash P^{S}(\hat{\psi}_{t}\odot\psi_{t})$
@@
    $\hat{\psi}_{T}\leftarrow(K_{T}\odot G_{T})^{\top}(\hat{\psi}_{T-1}\odot(\mathbf{1}_{n_{X}}\otimes u_{T-1}))$
@@
         $u_{T}\leftarrow\upmu_{T}\oslash P^{S}(\hat{\psi}_{t})$
@@
             $\psi_{t}\leftarrow(K_{t+1}\odot G_{t+1})(\psi_{t+1}\odot(\mathbf{1}_{n_{X}}\otimes u_{t+1}))$
@@
             $G_{t}\leftarrow\exp(\operatorname{diag}(\gamma_{t-1}^{(k+1)})\Delta/\varepsilon)$
@@
    $\psi_{0}\leftarrow(K_{1}\odot G_{1})(\psi_{1}\odot(\mathbf{1}_{n_{X}}\otimes u_{1}))$
@@
Storing the tensor $Q$ quickly becomes impossible as the number of marginals grows and consequently we cannot access the full transport $Q$ in practice. Fortunately, it is often enough to obtain a set of sub-transports of the form $P_{t_{1},t_{2}}(Q)$, where $t_{1},t_{2}\in[T]$ are such that $t_{1}<t_{2}$. For example, the corresponding problem value — the lower bound on the price — can be obtained without explicitly forming and storing the full tensor. See this by recalling that
$\langle\Phi,Q\rangle=\sum_{t\in[T]\backslash 0}\langle\Phi_{t},P_{t-1,t}(Q)\rangle,$
where the sub-transports $P_{t-1,t}(Q)$ for $t\in[T]\backslash 0$ are given by Corollary 4.4.
@@ 5 Computational results
In this section, we apply our computational framework to solve a number of MOT problems: First, we consider problems for which the optimal solution is known analytically; comparison of the computed solution with the known solutions allows for verifying that the method works as intended. We then conclude by solving a MOT problem where the optimal solution is not known.
@@ 5.1 Monotone transport
A well-known example from classical OT theory is the monotone transport plan, which is concentrated on the graph of an increasing function (see, for example, [60, p. 75]). It is also known as the Hoeffding-Fréchet coupling and solves the upper bound bi-marginal OT problem when the cost function $\phi:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ satisfies certain conditions [^5] The cost function $\phi$ should be upper semi-continuous and of linear growth. — in particular, the cross derivative $\partial_{s_{0}s_{1}}\phi(s_{0},s_{1})$ should exist and be strictly positive [36, Theorem 2.2]. Similarly, the anti-monotone transport plan is concentrated on the graph of a decreasing function; it solves the corresponding lower bound OT problem. There exists a martingale analogue to the monotone coupling; the left-monotone transport plan was introduced in [7, Definition 1.4] and shown to solve the bi-marginal MOT problem for two types of cost functions [7, Theorems 6.1 and 6.3]. It is concentrated on the graphs of two measurable functions and exhibits a V-shape when visualised in the plane. It was shown in [36, Theorem 5.1] that the left-monotone coupling solves the upper bound bi-marginal MOT problem, given that the derivative $\partial_{s_{0}s_{1}s_{1}}\phi(s_{0},s_{1})$ exists and is strictly positive. Since the monotone couplings are visually easily recognised, they serve as a clear first demonstrating example.
@@
Let $\phi(s_{0},s_{1})=(\log(s_{1}/s_{0}))^{2}$ and let the initial marginal $\mu_{0}$ be uniform with support on $n_{S_{0}}=600$ atoms, evenly distributed over the interval $[1.25,1.75]$. Similarly, let the terminal marginal $\mu_{1}$ be uniform with support on $n_{S_{1}}=1200$ atoms, evenly distributed over the interval $[1,2]$. It is easily verified that $\partial_{s_{0}s_{1}s_{1}}\phi>0$ and $\partial_{s_{0}s_{1}}\phi<0$ on the supports of $\mu_{0}$ and $\mu_{1}$ and that the marginals are in convex order; the optimal solutions for the corresponding upper-bound MOT and OT problems are thus the left-monotone and the anti-monotone couplings, respectively.
@@ 5.2 Late and early transports
$\phi(s_{0},\dots,s_{T})=\frac{1}{T+1}\sum_{t\in[T]}f(s_{t}),\quad s_{0},\dots s_{T}\in\mathcal{S},$ (38)
for some convex function $f:\mathbb{R}\rightarrow\mathbb{R}$, and where the initial and terminal marginals $\mu_{0}$ and $\mu_{T}$ are the only given marginals. As always, it is assumed that $\mu_{0}$ and $\mu_{T}$ are in convex order. Two particular martingale couplings are of interest here, the coupling that corresponds to not moving at all before $t=T-1$, then performing the full transport between $t=T-1$ and $t=T$, and the coupling that corresponds to performing the full transport between $t=0$ and $t=1$, then keeping the process constant until $t=T$. We refer to these couplings as late and early transport, respectively. The below result follows immediately from linearity of the expected value, since for any model $(\Omega,\mathcal{F},\mathbb{Q},S)$,
$\mathbb{E}_{\mathbb{Q}}\left[f(S_{T})\right]\geq\mathbb{E}_{\mathbb{Q}}\left[f(S_{T-1})\right]\geq\dots\geq\mathbb{E}_{\mathbb{Q}}\left[f(S_{1})\right]\geq\mathbb{E}_{\mathbb{Q}}\left[f(S_{0})\right],$
by convexity of $f$. See [57] for a similar result.
@@
$\phi_{1}(s_{0},s_{1})$ $=\frac{1}{T+1}\left(f(s_{0})+f(s_{1})\right),$ $s_{0},s_{1}\in\mathcal{S},$
$\phi_{t}(s_{t})$ $=\frac{1}{T+1}f(s_{t}),$ $s_{t}\in\mathcal{S},\quad t=2,\dots,T,$
@@
Figure 2: Marginal distributions corresponding to the optimal solutions of the upper- and lower bound MOTs from Example 5.3. The given initial marginal $\mu_{0}$ is displayed in light blue bars, while the given terminal marginal $\mu_{T}$ is shown in light red bars. The marginals $P_{t}(Q)$ corresponding to the computed transport plan $Q$ are shown for $t\in\{0,10,20,30,40,50\}$ with solid lines. Figure 2(a) corresponds to the lower bound martingale transport plan while Figure 2(b) corresponds to the upper bound martingale transport plan.
A subset of the marginal distributions of the computed optimal martingale transport plans are displayed in Figure 2, along with the given marginals $\mu_{0}$ and $\mu_{50}$. The lower bound solution (computed using $\varepsilon=0.0065$) is shown in Figure 2(a); since the computed intermediate marginals are close to the initial marginal this solution corresponds to a late transport, as predicted by Proposition 5.2. The upper bound solution (computed using $\varepsilon=0.02$) is shown in Figure 2(b); for this solution, we note that the intermediate marginals are close to the terminal marginal. It is therefore an example of early transport, as given by Proposition 5.2. For both solutions, the small deviations of the intermediate marginals from the initial or terminal marginal are due to the regularisation.

### v2

@@ 4.3 Summary of the full algorithm
We have now arrived at a form of the optimality conditions that allows for efficient computation of the optimal dual variables and we can therefore construct a coordinate dual ascent method by cyclically fixing all but one variable and optimise over the remaining variable by selecting it so that it satisfies the corresponding optimality condition. We then move on to the next variable until all variables have been optimised over; one such cycle defines one iteration in the method.
Each iteration starts with the update of the variables $u^{\lambda}$ corresponding to the marginal constraints (20c). The closed-form formula for doing so is given by Theorem 4.5 as
$u_{t}^{\lambda_{t}}\leftarrow m_{t}\oslash P^{\mathcal{S}}(\hat{\psi}_{t}\odot\psi_{t}),\quad t\in\mathcal{T}.$
From the above we note that is suffices to manipulate vectors in order to update $u^{\lambda}$, something that enables fast computation.
We then proceed to updating the dual variables $\gamma$ representing the martingale constraints (20b). This time, no closed-form formula exists for $\gamma$ — the optimality condition thus has to be solved numerically. Theorem 4.5 provides a form of the equations that allows for efficiently doing so by application of Newton’s method. In order to do so, define a family of functions $\alpha_{t}:\mathbb{R}^{n_{t-1}}\rightarrow\mathbb{R}^{n_{t-1}}$ for $t\in[T]\backslash 0$ via
$\alpha_{t}(z):=\begin{cases}\big{(}K_{t}\odot\exp(\operatorname{diag}(z)\Delta_{t}/\varepsilon)\odot\Delta_{t}\big{)}\big{(}\psi_{t}\odot(1_{n_{t}^{\mathcal{X}}}\otimes u_{t}^{\lambda_{t}})\big{)},&t\in\mathcal{T}\\ \big{(}K_{t}\odot\exp(\operatorname{diag}(z)\Delta_{t}/\varepsilon)\odot\Delta_{t}\big{)}\psi_{t},&t\notin\mathcal{T}.\end{cases}$
Then equation 25b can be written
$\begin{cases}\hat{\psi}_{t}\odot(1_{n_{t}^{\mathcal{X}}}\otimes u_{t}^{\lambda_{t}})\odot\alpha_{t+1}(\gamma_{t})=0_{n_{t}},&t\in[T-1]\cap\mathcal{T}\\ \hat{\psi}_{t}\odot\alpha_{t+1}(\gamma_{t})=0_{n_{t}},&t\in[T-1]\backslash\mathcal{T}.\end{cases}$
@@
$\gamma_{t}^{(k+1)}\leftarrow\gamma_{t}^{(k)}-\theta^{(k)}\odot J^{\alpha}_{t+1}(\gamma_{t}^{(k)})^{-1}\alpha_{t+1}(\gamma_{t}^{(k)}),\quad k=1,2,\dots,$ (33)
where $\theta^{(k)}\in\mathbb{R}^{n_{t}}$ is a vector containing the respective step lengths to be used for each variable, such that $\theta^{(k)}(j)\in[0,1]$ for $j\in\mathcal{I}_{t}$. It is determined by vectorised line search. In the above, $J^{\alpha}_{t+1}$ denotes the Jacobian matrix associated with $\alpha_{t+1}$; note that $\alpha_{t+1}$ is a vector-valued function whose $j^{\text{th}}$ component is independent of any other components of $\gamma_{t}$ than $\gamma_{t}(j)$, $j\in\mathcal{I}_{t}$. There are no cross dependencies. Therefore, the Jacobian is a diagonal matrix and given by $J^{\alpha}_{t}(\gamma_{t-1})=\operatorname{diag}(\beta_{t}(\gamma_{t-1}))$, where $\beta_{t}:\mathbb{R}^{n_{t-1}}\rightarrow\mathbb{R}^{n_{t-1}}$ for $t\in[T]\backslash 0$ is defined as
$\beta_{t}(z):=\begin{cases}\varepsilon^{-1}\big{(}K_{t}\odot\exp(\operatorname{diag}(z)\Delta_{t}/\varepsilon)\odot\Delta_{t}\odot\Delta_{t}\big{)}\big{(}\psi_{t}\odot(1_{n_{t}^{\mathcal{X}}}\otimes u_{t}^{\lambda_{t}})\big{)},&t\in\mathcal{T}\\ \varepsilon^{-1}\big{(}K_{t}\odot\exp(\operatorname{diag}(z)\Delta_{t}/\varepsilon)\odot\Delta_{t}\odot\Delta_{t}\big{)}\psi_{t},&t\notin\mathcal{T}.\end{cases}$
@@
$\gamma_{t}^{(k+1)}\leftarrow\gamma_{t}^{(k)}-\theta^{(k)}\odot\alpha_{t+1}(\gamma_{t}^{(k)})\oslash\beta_{t+1}(\gamma_{t}^{(k)}),\quad k=1,2,\dots$
The inversion of the Jacobian and the matrix-vector multiplication used in (33) are thus avoided and replaced by the elementwise division of one vector with another, allowing for efficient evaluation of the dual optimality condition (25b). Newton’s method is therefore particularly well suited for this step in the method.
Putting things together, we have completed the derivation of our algorithm — see Algorithm 2 for the full method. Convergence of the algorithm can be shown given that there, by Lemma 4.3, exists a feasible solution that is positive for all variables (cf. [36, Theorem 4.1 and Proposition 1]). Convergence of the coordinate dual ascent then follows from [51, Theorem 2 and Section 5].
Given: matrices $\{K_{t}\}_{t\in[T]\backslash 0}$ and $\{\Delta_{t}\}_{t\in[T]\backslash 0}$, index set $\mathcal{T}$, vectors $\{m_{t}\}_{t\in\mathcal{T}}$, scalar $\varepsilon>0$
Initialise: $u_{t}^{\lambda_{t}}\leftarrow 1_{n_{t}^{\mathcal{S}}}$ for $t\in[T]$
Initialise: $\gamma_{t}\leftarrow 1_{n_{t}}$ for $t\in[T-1]$
Initialise: $G^{\gamma_{t-1}}_{t}\leftarrow\exp(\operatorname{diag}(\gamma_{t-1})\Delta_{t}/\varepsilon)$ for $t\in[T]\backslash 0$
Initialise: $\hat{\psi}_{0}\leftarrow 1_{n_{0}}$
Initialise: $\psi_{T}\leftarrow 1_{n_{T}}$
Initialise: $\psi_{t}\leftarrow(K_{t+1}\odot G_{t+1})(\psi_{t+1}\odot(1_{n_{t+1}^{\mathcal{X}}}\otimes u_{t+1}^{\lambda_{t+1}}))$ for $t\in[T-1]$
@@
         $u^{\lambda_{0}}_{0}\leftarrow m_{0}\oslash P^{\mathcal{S}_{0}}(\psi_{0})$
@@
         $\hat{\psi}_{t}\leftarrow(K_{t}\odot G^{\gamma_{t-1}}_{t})^{\top}(\hat{\psi}_{t-1}\odot(1_{n_{t-1}^{\mathcal{X}}}\otimes u_{t-1}^{\lambda_{t-1}}))$
@@
             $u_{t}^{\lambda_{t}}\leftarrow m_{t}\oslash P^{\mathcal{S}_{t}}(\hat{\psi}_{t}\odot\psi_{t})$
@@
    $\hat{\psi}_{T}\leftarrow(K_{T}\odot G^{\gamma_{T-1}}_{T})^{\top}(\hat{\psi}_{T-1}\odot(1_{n_{T-1}^{\mathcal{X}}}\otimes u^{\lambda_{T-1}}_{T-1}))$
@@
         $u^{\lambda_{T}}_{T}\leftarrow m_{T}\oslash P^{\mathcal{S}_{T}}(\hat{\psi}_{T})$
@@
             $\psi_{t}\leftarrow(K_{t+1}\odot G^{\gamma_{t}}_{t+1})(\psi_{t+1}\odot(1_{n_{t+1}^{\mathcal{X}}}\otimes u_{t+1}^{\lambda_{t+1}}))$
@@
             $G^{\gamma_{t-1}}_{t}\leftarrow\exp(\operatorname{diag}(\gamma_{t-1}^{(k+1)})\Delta_{t}/\varepsilon)$
@@
    $\psi_{0}\leftarrow(K_{1}\odot G^{\gamma_{0}}_{1})(\psi_{1}\odot(1_{n_{1}^{\mathcal{X}}}\otimes u^{\lambda_{1}}_{1}))$
@@
Storing the tensor $\mathbf{Q}$ quickly becomes impossible as the number of marginals grows and consequently we cannot access the full transport $\mathbf{Q}$ in practice. Fortunately, it is often enough to obtain a set of sub-transports of the form $P_{t_{1},t_{2}}(\mathbf{Q})$, where $t_{1},t_{2}\in[T]$ are such that $t_{1}<t_{2}$. For example, the corresponding problem value — the lower bound on the price — can be obtained without explicitly forming and storing the full tensor. See this by recalling that $\langle\mathbf{\Phi},\mathbf{Q}\rangle=\sum_{t\in[T]\backslash 0}\langle\Phi_{t},P_{t-1,t}(\mathbf{Q})\rangle$, where the sub-transports $P_{t-1,t}(\mathbf{Q})$ for $t\in[T]\backslash 0$ are given by Corollary 4.8.
@@ 5 Computational results
In this section, we apply our computational framework to solve a number of MOT problems: First, we consider problems for which the optimal solution is known analytically; comparison of the computed solution with the known solution allows for verifying that the method works as intended. We then conclude by solving an MOT problem where the optimal solution is not known. In all examples the regularisation parameter $\varepsilon$ was chosen by trial and error to be as small as possible.
@@ 5.1 Monotone transport
A well-known example from classical OT theory is the monotone transport plan, which is concentrated on the graph of an increasing function (see, for example, [67, p. 75]). It is also known as the Hoeffding-Fréchet coupling and solves the upper bound bi-marginal OT problem when the cost function $\phi:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ satisfies certain conditions [^8] The cost function $\phi$ should be upper semi-continuous and of linear growth. — in particular, the cross derivative $\partial_{s_{0}s_{1}}\phi(s_{0},s_{1})$ should exist and be strictly positive [40, Theorem 2.2]. Similarly, the anti-monotone transport plan is concentrated on the graph of a decreasing function; it solves the corresponding lower bound OT problem. There exists a martingale analogue to the monotone coupling; the left-monotone transport plan was introduced in [7, Definition 1.4] and shown to solve the bi-marginal MOT problem for two types of cost functions [7, Theorems 6.1 and 6.3]. It is concentrated on the graphs of two measurable functions and exhibits a V-shape when visualised in the plane. It was shown in [40, Theorem 5.1] that the left-monotone coupling solves the upper bound bi-marginal MOT problem, given that the derivative $\partial_{s_{0}s_{1}s_{1}}\phi(s_{0},s_{1})$ exists and is strictly positive. Since the monotone couplings are visually easily recognised, they serve as a clear first demonstrating example.
@@
Let $\phi(s_{0},s_{1})=(\log(s_{1}/s_{0}))^{2}$ and let the initial marginal $\mu_{0}$ be uniform with support on $n_{0}^{\mathcal{S}}=600$ atoms, evenly distributed over the interval $[1.25,1.75]$. Similarly, let the terminal marginal $\mu_{1}$ be uniform with support on $n_{1}^{\mathcal{S}}=1200$ atoms, evenly distributed over the interval $[1,2]$. It is easily verified that $\partial_{s_{0}s_{1}s_{1}}\phi>0$ and $\partial_{s_{0}s_{1}}\phi<0$ on the supports of $\mu_{0}$ and $\mu_{1}$ and that the marginals are in convex order; the optimal solutions for the corresponding upper-bound MOT and OT problems are thus the left-monotone and the anti-monotone couplings, respectively.
@@ 5.2 Late and early transports
$\phi(s_{0},\dots,s_{T})=\frac{1}{T+1}\sum_{t\in[T]}f(s_{t}),\quad(s_{0},\dots s_{T})\in\mathcal{S}_{0}\times\dots\times\mathcal{S}_{T},$ (34)
for some convex function $f:\mathbb{R}\rightarrow\mathbb{R}$, and where the initial and terminal marginals $\mu_{0}$ and $\mu_{T}$ are the only given marginals. As always, it is assumed that $\mu_{0}$ and $\mu_{T}$ are in convex order. Two particular martingale couplings are of interest here, the coupling that corresponds to not moving at all before $t=T-1$, then performing the full transport between $t=T-1$ and $t=T$, and the coupling that corresponds to performing the full transport between $t=0$ and $t=1$, then keeping the process constant until $t=T$. We refer to these couplings as late and early transport, respectively. The below result follows immediately from linearity of the expected value, since for any model $(\Omega,\mathcal{F},\mathbb{Q},S)$ we have that $\mathbb{E}_{\mathbb{Q}}[f(S_{T})]\geq\mathbb{E}_{\mathbb{Q}}[f(S_{T-1})]\geq\dots\geq\mathbb{E}_{\mathbb{Q}}[f(S_{1})]\geq\mathbb{E}_{\mathbb{Q}}[f(S_{0})]$, by convexity of $f$. See [63] for a similar result.
@@
$\phi_{1}(s_{0},s_{1})$ $=\frac{1}{T+1}\left(f(s_{0})+f(s_{1})\right),$ $(s_{0},s_{1})\in\mathcal{S}_{0}\times\mathcal{S}_{1},$
$\phi_{t}(s_{t})$ $=\frac{1}{T+1}f(s_{t}),$ $s_{t}\in\mathcal{S}_{t},\quad t=2,\dots,T,$
@@
Figure 2: Marginal distributions corresponding to the optimal solutions of the upper- and lower bound MOTs from Example 5.3. The given initial marginal $\mu_{0}$ is displayed in light blue bars, while the given terminal marginal $\mu_{T}$ is shown in light red bars. The marginals corresponding to the computed transport plan are shown for $t\in\{0,10,20,30,40,50\}$ with solid lines. Figure 2(a) corresponds to the lower bound martingale transport plan while Figure 2(b) corresponds to the upper bound martingale transport plan.
A subset of the marginal distributions of the computed optimal martingale transport plans are displayed in Figure 2, along with the given marginals $\mu_{0}$ and $\mu_{50}$. The number of gridpoints used ranges from $n_{0}^{\mathcal{S}}=74$ to $n_{50}^{\mathcal{S}}=214$. The lower bound solution (computed using $\varepsilon=6.5\cdot 10^{-3}$) is shown in Figure 2(a); since the computed intermediate marginals are close to the initial marginal this solution corresponds to a late transport, as predicted by Proposition 5.2. The upper bound solution (computed using $\varepsilon=0.02$) is shown in Figure 2(b); for this solution, we note that the intermediate marginals are close to the terminal marginal. It is therefore an example of early transport, as given by Proposition 5.2. For both solutions, the small deviations of the intermediate marginals from the initial or terminal marginal are due to the regularisation.